OpenStudy (anonymous):

7a=4=-17 can you help me solve it?

OpenStudy (nenadmatematika):

7a=-17-4 7a=-21 a=-3 if you meant was 7a-4=-17

OpenStudy (anonymous):

sorry i meant to say 7a + 4 = -17

OpenStudy (precal):

if 7a+4=-17 subtract 4 to both sides 7a=-21 divide by 7 a=-3

OpenStudy (anonymous):

thank you